When performing a chi squaredχ2 test for independence in a contingency table with r rows and c​ columns, determine the​ upper-tail critical value of the test statistic in each of the following circumstances.

b. alpha = 0.01, r = 5, c = 3

c. alpha = 0.01, r = 5, c = 4

d. alpha = 0.01, r = 3, c = 4

e. alpha = 0.01, r = 4, c = 5

QUESTION: what are the critical values for b,c,d,e?
Answer #1

Part b

We are given

Degrees of freedom = df = (r – 1)*(c – 1) = (5 - 1)*(3 - 1) = 4*2 = 8

α = 0.01

Critical value = 20.09024

(by using Chi square table or excel)

Part c

We are given

Degrees of freedom = df = (r – 1)*(c – 1) = (5 - 1)*(4 - 1) = 4*3 = 12

α = 0.01

Critical value = 26.21697

(by using Chi square table or excel)

Part d

We are given

Degrees of freedom = df = (r – 1)*(c – 1) = (3 - 1)*(4 - 1) = 2*3 = 6

α = 0.01

Critical value = 16.81189

(by using Chi square table or excel)

Part e

We are given

Degrees of freedom = df = (r – 1)*(c – 1) = (4 - 1)*(5 - 1) = 3*4 = 12

α = 0.01

Critical value = 26.21697

(by using Chi square table or excel)